OpenStudy (anonymous):

Find the point on the terminal side of θ = 3π/4 that has a y coordinate of 1. Can somebody walk me through this? I don't really understand.

OpenStudy (shamim):

ur point is in the second coordinate

OpenStudy (shamim):

right?

OpenStudy (shamim):

3π/4=π/2+π/4

OpenStudy (anonymous):

OK, I think I partly solved the problem. Would you mind telling me if I am correct or not?

OpenStudy (shamim):

ok go on

OpenStudy (anonymous):

Using the unit circle, we know that the line on the second quadrant comes together with the line of the terminal side of θ to create a 45 degree angle. Am I correct so far?

OpenStudy (shamim):

u r right

OpenStudy (anonymous):

Next, since I know the length of the side of the triangle is 1 (y=1).

OpenStudy (anonymous):

I can use the tan function to get the other side. tan 45 = 1/x.

OpenStudy (shamim):

ya

OpenStudy (anonymous):

x is the adjacent side of θ that we are looking for.

OpenStudy (anonymous):

so x = 1?

OpenStudy (anonymous):

and that solved the problem, right?

OpenStudy (shamim):

u know ur point is in the second quadrunt

OpenStudy (shamim):

x=-1

OpenStudy (anonymous):

Oh, I am confused now. How would that get along with the equation?

OpenStudy (anonymous):

Like tan 45 = 1/x. How do I get x = -1 on that?

OpenStudy (anonymous):

Or I will just get x = 1 and then tell since it's in the second quadrant, x = -1?

OpenStudy (shamim):

u know all value of x in the second quadrunt in negative

OpenStudy (anonymous):

So I will go with "I will just get x = 1 and then tell since it's in the second quadrant, x = -1" ?

OpenStudy (shamim):

ys

OpenStudy (anonymous):

OK, I got it. Thank you so much for sticking with me. I appreciate it a lot!

OpenStudy (shamim):

i also enjoyed it